log5(x-1)=4+log5(x-2)
Answer by nerdybill(7384)

log5(x-1)=4+log5(x-2)
.
Begin by moving all the log5 terms to the left:
log5(x-1) - log5(x-2) = 4
log5[(x-1)/(x-2)] =4
(x-1)/(x-2) = 5^4
(x-1)/(x-2) = 625
(x-1) = 625(x-2)
x-1 = 625x - 1250
-624x-1 = -1250
-624x = -1249
x = 2.002